# On the Convergence Theory of Pipeline Gradient-based Analog In-memory Training

### v2

@@ I Introduction
Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications. As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive. For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [1]. In response to the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators has emerged. The ubiquitous fully connected and convolutional layers, whose core operation is matrix-vector multiplication (MVM), form the fundamental backbone of both training and inference as the demand for efficient processing intensifies. Typically, each forward and backward pass requires millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
@@
The promise of analog computing. To this end, a promising solution to accelerate the MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage. The DNN weights are stored in analog crossbar tiles consisting of resistive elements [2] where the conductance of the resistive element represents the weight magnitude. Unlike traditional Von Neumann accelerators such as GPUs and TPUs, which move weights from memory to the processor during computation, AIMC accelerators keep the weights stationary and perform MVM operations in memory. Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Fig. 1.
@@
The curse of data parallelism. While efficient computation is beneficial, in-memory computing imposes scalability limitations. Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training. In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different accelerators to support gradient descent with large batches [4, 5]. By contrast, in the AIMC architecture, trainable weights are stored in the analog tiles as the conductances of resistive elements; thus, copying weights is expensive and prone to error. Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
@@
Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [6]. Pipeline parallelism partitions a large model into a series of stages distributed across multiple accelerators, with each stage corresponding to the computation of a few layers. After an accelerator processes one datum (or batch), it passes the intermediate outputs to the next accelerator and begins processing the next datum. Pipeline parallelism overlaps computation across different accelerators, increasing system throughput.
@@
Synchronous and asynchronous parallelism. As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [7, 8], Deepspeed [9], FSDP [10], Megetron-LM [11], Colossal-AI [12], just to name a few. In their implementation, a gradient mini-batch (e.g., with batch size 128) is split into micro-batches (e.g., 8 micro-batches with batch size 16) and filled into the pipeline. The model weights are updated after all the micro-batches are processed. Featuring the synchronization operation at the end of each mini-batch, this pipeline paradigm is referred to as synchronous pipeline. However, during the fill and drain phases, the pipeline is not fully occupied, which limits the device utilization. To overcome the issue, asynchronous pipeline is proposed, which continuously fills the micro-batch into the pipeline and updates the weights once the gradient is available. Asynchronous pipeline increases the device utilization, while the weights for gradient computing can be stale, leading to a biased training dynamic.
@@
Despite two sources of bias, this paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and SGD on AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
@@ I-A Main results
This paper aims to understand the impact of bias and investigate the computational complexities and acceleration effect of Analog-SGD-AP. We highlight the unique technical challenges compared with the analysis of the asynchronous pipeline on digital accelerators as follows.
  1. T1)
    Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias. To be specific, consider an analog weight $W_{k}$ on a resistive crossbar array and a desired update $\Delta W_{k}$. The updated weight becomes
    $\textsf{Analog Update}\quad W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}^{(m)}$ (1)
    where $\tau$ is a hardware-specific constant, $|\,\cdot\,|$ is element-wise absolute value, and $\odot$ is the element-wise production. Unlike the accurate digital training, the last bias term in (1) is unique to analog accelerators. This bias can not be eliminated within a reasonable cost and leads to challenges to the analysis. In section II-B, we provide more background.
  2. T2)
    Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [13]. However, in the asynchronous pipeline, the forward pass is performed using stale weights, while the backward pass uses the latest weights. Consequently, their inconsistent multiplication used for weight updates is not a valid gradient, and this error accumulates during training. The stale signals in DNN render the update increment used in Analog-SGD-AP no longer a valid gradient, making its convergence behavior unclear.
  3. T3)
    Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis. Without considering a specific multi-layer architecture, prior work [14] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
@@ 3. T3)
TABLE I: Comparison of different pipeline strategies. Sample complexity is the number of samples needed to achieve a training loss $\leq\varepsilon$. Computing density is the average portion of active accelerators. In the table, $M$ and $B$ are the number of stages and micro-batches, respectively. Wall-clock complexity measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by computation density.
Overcoming the challenges, this paper answers the question above and provides both a convergence guarantee and computational complexities. The key contributions are as follows.
  1. C1)
    We consider the training of a multi-layer DNN, which consists of linear layers and non-linear activation functions. We formulate the training dynamics of Analog-SGD-AP, which reflects the stale weights in the asynchronous pipeline and the bias from the AIMC accelerators’ imperfection. We investigate the smoothness of the training objective as well as the robustness of the network against stale weights.
  2. C2)
@@
  3. C3)
    We verify the speedup of Analog-SGD-AP against Analog-SGD-WOP and Analog-SGD-SP in AIMC accelerator simulator, AIHWKIT. Numerical simulation results show that despite a slightly larger sample complexity, Analog-SGD-AP achieves the same accuracy with fewer clock cycles due to its high computation density, which is consistent with the theoretical guarantee. Furthermore, empirical results show that Analog-SGD-AP achieves linear speedup at least within the range of 1-8 accelerators.
@@ 3. C3)
To put our work in context, we review prior art that we group into the following two categories.
Gradient-based training on AIMC accelerators. Analog training has shown promising early successes in certain tasks [15, 16, 17], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware. However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [18]. Numerous algorithms have been proposed to address these issues. TT-v1 [19] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error. To demonstrate its efficiency, [20] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee. Building on that, [21] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods. To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [22] introduces a digital array. Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [23] are proposed to mitigate the impact of inaccurate symmetric point correction [24]. To overcome the limit update granularity issue, [25] proposed a multi-tile training paradigm. However, the mentioned works focus on training single AIMC accelerators. Instead, this work aims to scale up the system by parallelizing model computation.
Asynchronous pipeline. Existing works proposed methods to alleviate the stale weight issues in the asynchronous pipeline. For instance, [26] proposed weight-prediction approaches to compensate for the delay, PipeDream [27] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [28] and [29] propose only overlapping gradient computation and communication to ensure the delay is one. These algorithms are not for analog pipeline training. Researchers also attempted to design analog pipeline circuits to accelerate the training [30, 31], but they focus on the synchronous pipeline. On the theory side, there are few studies on pipeline training for digital accelerators such as [14, 32], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms. As [27] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
@@ II-A Training problem and back-propagation
$\textsf{Backward}\qquad\delta^{(m)}=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}=(\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$ (4)
@@
$\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}=\delta^{(m)}\otimes x^{(m)}$ (5)
@@
Rank-update Unlike digital accelerators, AIMC accelerators use rank-update to change the matrix, which leverages the outer product structure as shown in (5). To do so, independent random voltage pulses with probability proportional to the elements of $\delta^{(m)}$ and $x^{(m)}$ are applied to columns and rows, respectively. Consequently, the $(i,j)$-th element $W]_{i,j}$ receives a pulse coincident with probability $\propto[\delta^{(m)}]_{i}[x^{(m)}]_{j}$. Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, rank-update has only $O(BN)$ complexity while the update on digital accelerators has $O(BN^{2})$ complexity to update an $N\times N$ matrix. Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [15].
@@ II-B Non-ideal update on AIMC accelerator
Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [19, 20, 21], the update on AIMC accelerators exhibits asymmetric bias. To be specific, for any desired update $\Delta W_{k}^{(m)}$, the update on $W_{k}$ is scaled by response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$. Considering the weight $W_{k}$ at iteration $k$ and the expected update $\Delta W\in{\mathbb{R}}^{D}$, we express the asymmetric update as $W_{k+1}=U(W_{k},\Delta W)$ with the update function $U:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ defined by
@@
Defining the symmetric and asymmetric components as $F(w):=\frac{1}{2}(q_{-}(w)+q_{+}(w))$ and $G(w):=\frac{1}{2}(q_{-}(w)-q_{+}(w))$, we write the update in (6) as $U(w,\Delta w)=w+\Delta w\cdot F(w)-|\Delta w|\cdot G(w)$. This paper consider a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$ as suggested by [20], we reach the dynamics (1).
@@ II-C Synchronous and asynchronous pipelines
Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one accelerator and necessitates model parallelism. In model parallelism, the forward and backward computations are partitioned into a sequence of stages, which, in general, consist of a consecutive set of layers in the model. Each stage is mapped to one digital or analog accelerator. Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$. Without loss of generality, this paper assumes that one stage contains only one layer and the micro-batch size $B_{\text{micro}}=1$. Under this assumption, $\xi_{k,b}$ denotes the $b$-th micro-batch which contains only the $b$-th data. Consequently, the number of micro-batches is the mini-batch size, i.e., $B=B_{\text{mini}}$. This paper treats a stage as the minimal scheduling unit and investigates the time complexity across different scheduling strategies. To compare time costs, we define the basic time unit, clock cycle, as the minimum time required for all devices to process one micro-batch. Based on scheduling strategies, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining. Vanilla model parallelism without pipeline. Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data sample, as previously mentioned) for iteration $k$. Vanilla model parallelism processes micro-batches sequentially. In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$ according to (3), while in the backward pass stage $m$ achieves the backward signal $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage according to (4). The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stashed when the gradient for the $b$-th micro-batch is completed. After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method. Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}^{(m)}$, which iterates over $b\in[B_{\text{mini}}]$ by
@@
Notice that for digital SGD, the last term in the RHS (RHS) of (7) disappears, and hence (7) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$. To fully release the potential of rank-update, AIMC accelerators follow (7) to update its weight, which in turn complicates the convergence analysis.
@@
Synchronous pipeline. Vanilla model parallelism is simple to implement. However, most devices remain idle, reducing system throughput. Existing work has already attempted to design an architecture that supports the pipeline training [30, 31]. They faithfully implement the iteration (7) but overlap the computing inside a mini-batch. Featuring synchronization at the end of a mini-batch, this stage schedule mechanism is called synchronous pipeline. In the vanilla non-pipeline version, stage $m$ remains idle after completing the forward computation of the $b$-th micro-batch until the backward signal returns from stage $m+1$. In the synchronous pipelining, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any. Similar to the vanilla model parallelism, the weights will be updated after processing all micro-batches. Consequently, for each mini-batch, most of the devices need to wait until the last micro-batch is processed. After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (7). To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch. During the pipeline filling and draining periods, parts of the devices are idle; see the middle of Fig. 2
@@
Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k},{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$. (9)
Imperfect gradient-based training on AIMC accelerators. After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights. However, there is a bias due to the hardware imperfection [20]. Replacing $\Delta W_{k}$ with $\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ in (1), where $\alpha>0$ is a positive learning rate, we reach the update dynamics of weights

### v3

@@ I Introduction
Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications. As the model sizes and dataset volumes grow exponentially, DNN training becomes time-consuming and resource-intensive. For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [touvron2023llama2]. In response to the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators has emerged. The ubiquitous fully connected and convolutional layers, whose core operation is matrix-vector multiplication (MVM), form the fundamental backbone of both training and inference as the demand for efficient processing intensifies. Typically, each forward and backward pass requires millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
@@
The promise of analog computing. To this end, a promising solution to accelerate MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage. The DNN weights are stored in analog crossbar tiles consisting of resistive elements [jain2019neural] where the conductance of the resistive element represents the weight magnitude. Unlike traditional Von Neumann accelerators such as GPUs and TPUs, which move weights from memory to the processor during computation, AIMC accelerators keep the weights stationary and perform MVM operations in memory. Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Fig. 1.
@@
The curse of data parallelism. While efficient computation is beneficial, in-memory computing imposes scalability limitations. Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training. In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [li2020pytorch], where DNN weights are replicated to different accelerators to support gradient descent with large batches [goyal2017accurate, you2020Large]. By contrast, in the AIMC architecture, trainable weights are stored in the analog tiles as the conductances of resistive elements; thus, copying weights is expensive and prone to error. Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
@@
Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [xu2021parallelizing]. Pipeline parallelism partitions a large model into a series of stages distributed across multiple accelerators, with each stage corresponding to the computation of a few layers. After an accelerator processes one datum (or batch), it passes the intermediate outputs to the next accelerator and begins processing the next datum. Pipeline parallelism overlaps computation across different accelerators, increasing system throughput.
@@
Synchronous and asynchronous parallelism. As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [huang2019gpipe, kim2020torchgpipe], Deepspeed [ren2021zero], FSDP [zhao2023pytorch], Megetron-LM [shoeybi2019megatron], Colossal-AI [li2023colossal], just to name a few. In their implementation, a gradient mini-batch (e.g., with batch size 128) is split into micro-batches (e.g., 8 micro-batches with batch size 16) and filled into the pipeline. The model weights are updated after all the micro-batches are processed. Featuring the synchronization operation at the end of each mini-batch, this pipeline paradigm is referred to as synchronous pipeline. However, during the fill and drain phases, the pipeline is not fully occupied, which limits the device utilization. To overcome the issue, asynchronous pipeline is proposed, which continuously fills the micro-batch into the pipeline and updates the weights once the gradient is available. Asynchronous pipeline increases the device utilization, while the weights for gradient computing can be stale, leading to a biased training dynamic.
@@
This paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
@@ I-A Main results
This paper aims to understand the impact of bias and investigate the computational complexities and acceleration effect of Analog-SGD-AP. We highlight the unique technical challenges compared with the analysis of the asynchronous pipeline on digital accelerators as follows. T1) Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias. To be specific, consider an analog weight $W_{k}$ on a resistive crossbar array and a desired update $\Delta W_{k}$. The updated weight becomes
$W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}$ (1)
where $\tau$ is a hardware-specific constant, $|\,\cdot\,|$ is element-wise absolute value, and $\odot$ is the element-wise production. Unlike digital training, the last term in (1) is unique to analog accelerators. This bias can not be eliminated within a reasonable cost and leads to challenges to the analysis. In section II-B, we provide more background. T2) Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [sun2017asynchronous]. However, in the asynchronous pipeline, the forward pass is performed using stale weights, while the backward pass uses the latest weights. Consequently, their inconsistent multiplication used for weight updates is not a valid gradient, and this error accumulates during training. The stale signals in DNN render the update increment used in Analog-SGD-AP no longer a valid gradient, making its convergence behavior unclear. T3) Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis. Without considering a specific multi-layer architecture, prior work [huo2018decoupled] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
@@
TABLE I: Comparison of different pipeline strategies. Sample complexity is the number of samples needed to achieve a training loss $\leq E+\varepsilon$, where $E$ is an inevitable asymptotic error due to hardware imperfections. Computing density is the average portion of active accelerators. In the table, $M$ and $B$ are the number of stages and micro-batches, respectively. Wall-clock complexity measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by computation density. See Theorems 1 and 2 for iteration complexity analysis; and Corollary 1 and 2 for sample and wall-clock complexity analysis.
@@
Prior art can be grouped into two categories.
Gradient-based training on AIMC accelerators. Analog SGD implemented by pulse update is extremely efficient [gokmen2016acceleration, yao2017face, wang2019situ], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware. However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [burr2015experimental]. Numerous algorithms have been proposed proposed to address these issues. TT-v1 [gokmen2020] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error. To demonstrate its efficiency, [wu2024towards] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee. Building on that, [wu2025analog] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods. To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [gokmen2021] introduces a digital array. Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [rasch2024fast] are proposed to mitigate the impact of inaccurate symmetric point correction [onen2022neural]. To overcome the limit update granularity issue, [li2025memory] proposed a multi-tile training paradigm. However, the mentioned works focus on training single AIMC accelerators. Instead, this work aims to scale up the system by parallelizing model computation.
Asynchronous pipeline. Existing works proposed methods to alleviate the stale weight issues in the asynchronous pipeline. For instance, [kosson2021pipelined] proposed weight-prediction approaches to compensate for the delay, PipeDream [narayanan2019pipedream] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [chen2022sapipe] and [li2018pipe] propose only overlapping gradient computation and communication to ensure the delay is one. These algorithms are not for analog pipeline training. Researchers also attempted to design analog pipeline circuits to accelerate the training [shafiee2016isaac, song2017pipelayer], but they focus on the synchronous pipeline. On the theory side, there are few studies on pipeline training for digital accelerators such as [huo2018decoupled, xu2020acceleration], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms. As [narayanan2019pipedream] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
@@ II-A Training problem and back-propagation
$\textsf{Backward}\qquad\delta^{(m)}=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m+1)}}\frac{\partial z^{(m+1)}}{\partial x^{(m+1)}}\frac{\partial x^{(m+1)}}{\partial z^{(m)}}$ (4)
$=$ $\ (\delta^{(m+1)}W^{(m)})\nabla g^{(m)}(z^{(m)}).$
@@
$\textsf{Gradient}\qquad\nabla_{W^{(m)}}f(W;\xi)=$ $\ \frac{\partial f(W;\xi)}{\partial z^{(m)}}\frac{\partial z^{(m)}}{\partial W^{(m)}}$ (5)
$=$ $\ \delta^{(m)}\otimes x^{(m)}$
@@
Rank-update Unlike digital accelerators, AIMC accelerators use rank-update to change the matrix, which leverages the outer product structure as shown in (5). To do so, independent random voltage pulses with probability proportional to the elements of $\delta^{(m)}$ and $x^{(m)}$ are applied to columns and rows, respectively. Consequently, the $(i,j)$-th element $W]_{i,j}$ receives a pulse coincident with probability $\propto[\delta^{(m)}]_{i}[x^{(m)}]_{j}$. Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, rank-update has only $O(BN)$ complexity while the update on digital accelerators has $O(BN^{2})$ complexity to update an $N\times N$ matrix. Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [gokmen2016acceleration].
@@ II-B Non-ideal update on AIMC accelerator
Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [gokmen2020, wu2024towards, wu2025analog], the update on AIMC accelerators exhibits asymmetric bias. To be specific, for any desired update $\Delta W_{k}^{(m)}$, the update on $W_{k}$ is scaled by response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$. Considering the weight $W_{k}$ at iteration $k$ and the expected update $\Delta W\in{\mathbb{R}}^{D}$, we express the asymmetric update as $W_{k+1}=U(W_{k},\Delta W)$ with the update function $U:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ defined by
@@
Defining the symmetric and asymmetric components as $F(w):=\frac{1}{2}(q_{-}(w)+q_{+}(w))$ and $G(w):=\frac{1}{2}(q_{-}(w)-q_{+}(w))$, we write the update in (6) as $U(w,\Delta w)=w+\Delta w\cdot F(w)-|\Delta w|\cdot G(w)$. This paper considers a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$, which leads to the dynamics (1). The linear forms of $q_{+}(w)$ and $q_{-}(w)$ can be viewed as a first-order approximation of more general nonlinear device response functions along the training trajectory. In this model, $\tau$ controls how strongly the update magnitude depends on the current conductance state. Physically, a larger $\tau$ means weaker state dependence (more symmetric updates), while a smaller $\tau$ means stronger asymmetry and stronger bias during training. Therefore, $\tau$ is a hardware-specific constant. Refer to [wu2024towards] for more details about the background of this model.
Remark 1.
As established in [gokmen2020, wu2024towards], the asymmetric update is one of the most significant challenges in analog training, and our theoretical framework explicitly captures this dominant source of non-ideality through the bias term in equation (1). In the theoretical analysis, we focus on the asymmetric update as the primary source of hardware imperfection to balance between faithfulness to the hardware and analytical tractability. Beyond the asymmetric update, the simulations additionally incorporate a comprehensive set of hardware imperfections to more accurately reflect real device behavior; see Section IV.
Beyond the update model in (6), additional hardware effects (e.g., output noise) can be viewed as bounded perturbations to the idealized update dynamics. Under such perturbations, the dominant convergence trend is expected to be preserved, while constants and the asymptotic error may increase by perturbation-dependent terms. A full study that explicitly incorporates these non-idealities is left for future work.
@@ II-C Synchronous and asynchronous pipelines
Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one accelerator and necessitates model parallelism. In model parallelism, the forward and backward computations are partitioned into a sequence of stages, which, in general, consist of a consecutive set of layers in the model. Each stage is mapped to one digital or analog accelerator. Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$. Without loss of generality, this paper assumes that one stage contains only one layer and the micro-batch size $B_{\text{micro}}=1$. Under this assumption, $\xi_{k,b}$ denotes the $b$-th micro-batch which contains only the $b$-th data. Consequently, the number of micro-batches is the mini-batch size, i.e., $B=B_{\text{mini}}$. This paper treats a stage as the minimal scheduling unit and investigates the time complexity across different scheduling strategies. To compare time costs, we define the basic time unit, clock cycle, which is the minimum time required for all devices to process one micro-batch. Note that this definition is based on the assumption that per-stage computation loads are balanced across all devices, so that all stages complete each micro-batch in approximately the same amount of time, which can be achieved via balanced partitioning [huang2019gpipe, narayanan2019pipedream, shoeybi2019megatron]. Based on scheduling strategies, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining. Vanilla model parallelism without pipeline. Let $\xi_{k,b},b\in[B]$ denote the $b$-th micro-batch (containing only one data sample, as previously mentioned) for iteration $k$. Vanilla model parallelism processes micro-batches sequentially. In the forward pass, stage $m$ achieves the output of the previews stage $x^{(m)}_{k,b}$ as input and outputs $x^{(m+1)}_{k,b}$ according to (3), while in the backward pass stage $m$ achieves the backward signal $\delta^{(m)}_{k,b}$ from next stage and sends the error $\delta^{(m-1)}_{k,b}$ to its previews stage according to (4). The signals $x^{(m)}_{k,b}$ and $\delta^{(m)}_{k,b}$ are stashed when the gradient for the $b$-th micro-batch is completed. After achieving all $\{(x^{(m)}_{k,b},\delta^{(m)}_{k,b}):b\in[B_{\text{mini}}]\}$, the weights will be updated using the gradient-based method. Beginning from $W_{k,0}=W_{k}$, Analog SGD update the weight as $W_{k+1}^{(m)}=W_{k,B}^{(m)}$, which iterates over $b\in[B_{\text{mini}}]$ by
$W_{k,b+1}^{(m)}=$ $\ W_{k,b}^{(m)}-\frac{\alpha}{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$
$\ -\frac{\alpha}{\tau B}|\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})|\odot W_{k,b}^{(m)}.$ (7)
@@
Notice that for digital SGD, the last term in the RHS (RHS) of (II-C) disappears, and hence (II-C) reduces to mini-batch SGD, i.e., $W_{k+1}^{(m)}=W_{k}^{(m)}-\frac{\alpha}{B}\sum_{b=1}^{B}\nabla_{W^{(m)}}f(W_{k};\xi_{k,b})$. To fully release the potential of rank-update, AIMC accelerators follow (II-C) to update its weight, which in turn complicates the convergence analysis.
@@
Synchronous pipeline. Vanilla model parallelism is simple to implement. However, most devices remain idle, reducing system throughput. Existing work has already attempted to design an architecture that supports the pipeline training [shafiee2016isaac, song2017pipelayer]. They faithfully implement the iteration (II-C) but overlap the computing inside a mini-batch. Featuring synchronization at the end of a mini-batch, this stage schedule mechanism is called synchronous pipeline. In the vanilla non-pipeline version, stage $m$ remains idle after completing the forward computation of the $b$-th micro-batch until the backward signal returns from stage $m+1$. In the synchronous pipelining, each stage sends the output signal $x^{(m+1)}_{k,b}$ to the next stage, while simultaneously starting to process the next forward signal $x^{(m)}_{k,b+1}$ or backward signal $\delta^{(m)}_{k,b-(M-m)}$, if any. Similar to the vanilla model parallelism, the weights will be updated after processing all micro-batches. Consequently, for each mini-batch, most of the devices need to wait until the last micro-batch is processed. After that, each device collects the gradients of the whole mini-batch and updates weights iteratively following (II-C). To implement that, a synchronization operation is required to wait for the completion of gradient computation before starting the next mini-batch. During the pipeline filling and draining periods, parts of the devices are idle; see the middle of Fig. 2
@@
Forward $\ {\tilde{x}}^{(0)}_{k}=x_{k},~~{\tilde{x}}^{(m+1)}_{k}=g^{(m)}({\tilde{z}}^{(m)}_{k})$, (8) ${\tilde{z}}^{(m)}_{k}=W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}$,~~ Backward $\ \tilde{\delta}^{(m)}_{k}=\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})$. (9)
Imperfect gradient-based training on AIMC accelerators. After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights. However, there is a bias due to the hardware imperfection [wu2024towards]. Replacing $\Delta W_{k}$ with $\alpha\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}$ in (1), where $\alpha>0$ is a positive learning rate, we reach the update dynamics of weights